OpenStudy (anonymous):

What's the equation that represents the line passing through the point (-3, -1)with a slope of 4.

OpenStudy (aroub):

y=mx+b , m=4 & (-3,-1) -1=4(-3)+b -1=-12+b 11=b =>y=4x+11